A popular news agency covering a local election will declare a winner when their sample data provides enough evidence that the proportion of all voters who will vote for that candidate is a majority—more than 50%.



If data from a random sample of 120 voters indicates that 66 of them are voting for a certain candidate, what value of z is the test statistic? (Round your answer to two decimal places.)
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.50

Ha : p > 0.50

= x / n = 66/120=0.55

P0 = 0.50

1 - P0 = 1-0.50=0.50

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.55-0.50/ [(0.50*0.50) /120 ]

= 1.09529

z =1.10